OpenStudy (howard-wolowitz):

If an object is dropped from a height, its downward speed theoretically increases linearly over time because the object is subject to the steady pull of gravity. Here are observational data on the speed of a ball dropped from a certain height at time x = 0: Time (seconds) X 0 0.2 0.4 0.6 0.8 Speed (m/sec) Y 0 1.92 3.58 6.01 7.88 11. Determine the least-squares regression line (LSRL). Write the equation below

OpenStudy (xmissalycatx):

Here's your graph first of all..

OpenStudy (xmissalycatx):

I might need @jigglypuff314 to help me. D:

OpenStudy (xmissalycatx):

or @ParthKohli

OpenStudy (howard-wolowitz):

isnt this my graph

OpenStudy (michele_laino):

please you don't draw the segments between experimental points. In other words, what we get is a scatter plot of @xMissAlyCatx

OpenStudy (michele_laino):

@Howard-Wolowitz

OpenStudy (michele_laino):

according to the scatter plot of @xMissAlyCatx we see that the relation between space and time is not a linear relation, since we can not fit such data with a straight line

OpenStudy (howard-wolowitz):

ok so my plot is correct for a scatterplot

OpenStudy (michele_laino):

yes! Nevertheless you have to delete the segments between experimental points

OpenStudy (howard-wolowitz):

how on I dont know what you mean? can you show me

OpenStudy (michele_laino):

Sorry, linear relation can be applied to your experimental data

OpenStudy (howard-wolowitz):

ok so you do believe that my scatter plot is 100% correct for the points given?

OpenStudy (michele_laino):

yes! Here is what I meant:

OpenStudy (howard-wolowitz):

ok, but thats what i have? so i dont understand

OpenStudy (michele_laino):

now, we have to conjecture this relation: \[\huge v = At + B\] where \(v\) is the speed after \(t\) seconds. Therefore we have to apply the usual formulas in order to get the values of both constants \(A,\;B\) starting from the coordinates of the experimental points

OpenStudy (howard-wolowitz):

are you talking about this question now: 11. Determine the least-squares regression line (LSRL). Write the equation below

OpenStudy (michele_laino):

yes!

OpenStudy (howard-wolowitz):

ok so i am with yoy so far with what you put

OpenStudy (howard-wolowitz):

how do we conjecture that relation

OpenStudy (michele_laino):

as I said before, we have to evaluate, starting from your experimental data, both constants \(A,\;B\)

OpenStudy (michele_laino):

here are the corresponding formulas:

OpenStudy (michele_laino):

please wait a moment I'm writing such formulas...

OpenStudy (michele_laino):

\[\Large \begin{gathered} \Delta = N\left( {\sum {t_i^2} } \right) - {\left( {\sum {{t_i}} } \right)^2} \hfill \\ \hfill \\ B = \frac{{\left( {\sum {t_i^2} } \right)\left( {\sum {{v_i}} } \right) - \left( {\sum {{t_i}} } \right)\left( {\sum {{t_i}{v_i}} } \right)}}{\Delta }, \hfill \\ \hfill \\ A = \frac{{N\left( {\sum {{t_i}{v_i}} } \right) - \left( {\sum {{t_i}} } \right)\left( {\sum {{v_i}} } \right)}}{\Delta } \hfill \\ \end{gathered} \]

OpenStudy (michele_laino):

where \(\Large N=5\)

OpenStudy (michele_laino):

of course, the index \(\Large i\) runs from \(\Large 1\) to \(\Large 5\)

OpenStudy (howard-wolowitz):

ok so we have to use the foeumla to get the final equation

OpenStudy (michele_laino):

that's right! Once we know both constants \(A,\;B\) we can write the above formula

OpenStudy (howard-wolowitz):

ok, the only part i dont get is how do we know them

OpenStudy (michele_laino):

what is the value for \(\Large \Delta\) ?

OpenStudy (howard-wolowitz):

thats the variable

OpenStudy (michele_laino):

Please wait, I'm checking your values...

OpenStudy (michele_laino):

I got: \(A=9.925\)

OpenStudy (michele_laino):

and \(B=-0.092\)

OpenStudy (howard-wolowitz):

show me your work please

OpenStudy (michele_laino):

ok! Please wait...

OpenStudy (michele_laino):

\[\Large \begin{gathered} \Delta = N\left( {\sum {t_i^2} } \right) - {\left( {\sum {{t_i}} } \right)^2} = 5 \cdot \frac{6}{5} - {2^2} = 2 \hfill \\ \hfill \\ B = \frac{{\left( {\sum {t_i^2} } \right)\left( {\sum {{v_i}} } \right) - \left( {\sum {{t_i}} } \right)\left( {\sum {{t_i}{v_i}} } \right)}}{\Delta } = \hfill \\ \hfill \\ = \frac{{\left( {6/5} \right) \cdot \left( {19.39} \right) - \left( 2 \right) \cdot \left( {11.726} \right)}}{2} = - 0.092, \hfill \\ \hfill \\ A = \frac{{N\left( {\sum {{t_i}{v_i}} } \right) - \left( {\sum {{t_i}} } \right)\left( {\sum {{v_i}} } \right)}}{\Delta } = \hfill \\ \hfill \\ = \frac{{5 \cdot \left( {11.726} \right) - \left( 2 \right) \cdot \left( {19.39} \right)}}{2} = 9.925 \hfill \\ \end{gathered} \]

OpenStudy (howard-wolowitz):

ok i have a question, I dont get what the question means by the equation... do we set up a new equation since the we have the values? thats confusing me sorry

OpenStudy (michele_laino):

we can say that the straight line of best fit, has the subsequent equation: \[\huge v\left( t \right) = - 0.092 + 9.925t\]

OpenStudy (michele_laino):

please try to draw such line, inside your scatter plot

OpenStudy (howard-wolowitz):

ok give me a second plz

OpenStudy (howard-wolowitz):



OpenStudy (michele_laino):

please, I don't see the straight line

OpenStudy (michele_laino):

for example, at \(t=0\) we have: \(v=-0.092\) so the first point is: \(P=(0,-0.092)\) whereas at \(t=1\), we have \(v=9.833)\), so the second point is: \(Q=(1,9.833)\) So, please try to draw a line which passes at point P and at point Q

OpenStudy (howard-wolowitz):

its n this correct?

OpenStudy (michele_laino):

no, I'm sorry, it is not correct

OpenStudy (howard-wolowitz):

no i mean i add the points on this correct?

OpenStudy (michele_laino):

please add your experimental points, to this graph:

OpenStudy (michele_laino):

second step: the second inequality, can be rewritten as below: \[\huge y \geqslant \frac{2}{3}x - 4\] am I right?

OpenStudy (michele_laino):

oops.. sorry I was answering to another student

OpenStudy (howard-wolowitz):

lol i wsa like im so confused

OpenStudy (howard-wolowitz):

ok so why would i put the points on your graph, your points arent accurate

OpenStudy (michele_laino):

so you can see that the best fit line, really fits to your experimental points

OpenStudy (michele_laino):

nevertheless, my equation above, solves completely your exercise

OpenStudy (howard-wolowitz):

i would like to see your final line plz